The Atomic Force microscope has been moved to the chemistry department. The group of Igmar Swart who works on the chemistry and physics of 2D materials (www) is dedicated to AFM imaging assuring frequent operation and service of the system. Contact Igmar, Joost van der Lit or Bart Gadella if you have imaging questions for this microscope.
Last Thursday, the 30th of October, the PC of the Leica SPE-II confocal crashed, this is already the second crash of the PC this year (we had the HD replaced just a half a year ago). We are trying to get the system running again in the next few days, but rely on ICT support and Leica service.
Data is secured and can be accessed when we plug the HD into another system, please ask the CCI staff to help getting data if you do not have a backup and really need it the next few days.
We bought a new shutter to rapidly block halogen light in time recordings using the EM-CCD camera of the NIKON imaging station. Previously, intervals between fluorescent observation and DIC recording required warming up the halogen lamp for at least 4 seconds. Now the light path is rapidly switchable in milliseconds and is incorporated into the NIS software and its Optical Configurations.
This week (Tue 8/7/2014) the hard disk of the SPEII system crashed. We had to remove it and re-install windows. The data on the extracted disk is very poorly readable and thus data and settings of users are essentially lost. Since all users are highly disciplined in their data backup procedures (?!) we can easily retrieve them. If not, contact us and we will help making new settings.
We recommend not to use data files older than 5 months for recovering acquisition settings, the updates of Leica software are not fully backward compatible and we have seen errors because of the older setting files. Best is to make new setting files from scratch in these cases.
On April 2nd, a Leica service engineer will try and fix the 561 nm laserline that very slowly starts. Also, the galvo x- and y- motors are going to be checked, they make strange sounds when stopping a scan, a recurring defect it seems.
Today, NIKON replaced the DIC NII-slider for 100x for the DIC-R variant, it solves the issue of spherical aberration caused by the DIC-II which was obvious in calibration bead imaging. The condensor DIC-R module works fine for the 60x lens with DIC NII slider in place.
We borrowed a stereoscope from the Biochemistry and Cell Biology department. It is “Little Red Riding Hood” and is very useful to prepare large specimen such as Oocytes, Zebrafish and C.elegans for microscopy.
Its a loan and can be taken back to B&C for teaching purposes, so we need to keep it clean and in perfect working order.
The NIKON system has been updated again, a 2 TB hard drive is installed that allows temporary storage untill you verified dataintegrity . We will imply an automated sync-system to move data from the C:\ drive to the new F:\ drive when the system is not in use.
A new cy5-cube has been placed in the 3rd A1R filter wheel set to allow sensitive imaging of far red dyes in the GaAsP detector, and outperforms the old filtercube several fold.
Users need to purchase their own stocks of imaging chambers, we were lucky to obtain a set of chambers from IBIDI and they perform well on our systems. We also have good experience with WPI 3 cm Fluodishes (having a 8 mm or 23 mm imaging well), initial problems with Labtech chambered coverglas slides are solved. Low cost solutions using ATTOfluor cell chambers clamping 25 mm round cs from Marienfeld are also available. It should be noted that mounting cs sometimes induces tension in the clamped cs which is released in an unpredictable manner, resulting in a loss of focus. Remember that objectives of high end imaging systems are calibrated to work with #1.5 (or 0.17 mm) thick glass coverslips.
Lastly, the TOKAI hit heater setup will be updated to allow better fluid replacement during imaging in the coming weeks.
dd October 2013
The A1R/Storm system is running satisfactory, we keep testing its capacities. NIS software has been updated to the latest 4.2 version with quite some improvements. Also the SPE2 has been serviced and software was updated. Please contact us if you have difficulties.
dd June 2013
The NIKON A1R/STORM system is running. A1R CLSM live cell imaging is performing quite well, we await filter changes to optimally image gfp / rfp variants. The STORM is still in its testing phase, especially dual/triple color imaging is challenging. The AFM has moved to the CCI and is operational, contact Bart Gadella for further info.
dd April 2013
The NIKON A1R/STORM system has been installed and the first imaging validations are being preformed. Please ask staff if you are interested to image on the machine. We will evaluate the time schedule and need for you to do your experiment on the machine.
dd March 2013:
A new state of the art microscope will be installed, the NIKON A1R/STORM system. We will be trained in the coming weeks on the A1R CLSM, the TIRF and STORM options.
The AFM from the B&C department will be moved to the CCI.
dd Jan 2013:
new Leica cryostate 3050S replaced the broken IRAS cryostate, thanks to collaborive effort of IRAS, DWM,Pathobiologie, I&I and GZD. Thank you all on behalf of the users for the continuation of this functionality at the CCI.
The OME image database on a LINUX server (with kind help of Aart deGroot) is working fine. We are exploring ways to properly implement it into the network to have central collection of the microscope data. We aim to store these for about half a year so users can acces the files easily and arrange definite storage capacity of their own.